OpenStudy (anonymous):

How many neutrons are in ? 74 32 106 42

OpenStudy (rushwr):

in which element ! I think part of the question is not there ~

OpenStudy (anonymous):

\[\frac{ 14 }{ 6 } C\]

OpenStudy (anonymous):

No wrong one sorry hang on

OpenStudy (rushwr):

neutrons in that is gonna be 8

OpenStudy (anonymous):

\[\frac{ 74 }{ 32 }Ge\]

OpenStudy (rushwr):

that would be 42 !

OpenStudy (rushwr):

32 is the no. of protons ( that is equal to the no. of electrons as well) 74 is the total summation of no. of neutrons and protons. Since we know that the no. of protons are 32. So the remaining from the 74 would be the neutrons. 74-32=42